# sample.bib

### v2

@@ 4.2 Utilization of LLMs in radiation oncology physics training and education
With the excellent performance of LLMs in answering radiation oncology physics questions evaluated in this study, it offers a promising avenue for their utilization in radiation oncology physics training and education to enhance learning efficiency and accessibility. For example, LLMs can assist in educating graduate students, postdocs, and medical physics residents as they prepare for challenging medical physics certification exams, such as the American Board of Radiology certification exams. LLMs have the unique advantage of providing immediate and context-specific explanations of medical physics concepts and are efficient at retrieving information from a wide range of topics. In busy clinical environments, where mentors may sometimes be occupied with clinical duties, LLMs can serve as a useful tool to address easy questions and have learners pay more attention to more complex problems. This leads to more focused and productive discussions with mentors. In summary, LLMs can bridge gaps in traditional learning settings and support self-directed study in radiation oncology physics.
@@ 5 Conclusion
We evaluated recently released LLMs using a method that randomly shuffled the answer options of radiation oncology physics questions. Our results demonstrated that these models achieved expert-level performance on these questions, with some even surpassing human experts with a majority vote. However, when the correct answer options were replaced with ”None of the above,” all models exhibited a steep decline in performance, suggesting room for improvement. Employing the technique of explaining-first and step-by-step instruction enhanced the reasoning capabilities of LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet. The performance of these recently released LLMs exhibited great potential for assisting in radiation oncology physics training and education.

### v3

@@ 4.2 Utilization of LLMs in radiation oncology physics training and education
With the excellent performance of LLMs in radiation oncology physics evaluated in this study, it offers a promising avenue for their utilization in radiation oncology physics education to enhance learning efficiency and accessibility. For example, LLMs can assist in educating graduate students, postdocs, and medical physics residents as they prepare for challenging medical physics certification exams, such as the American Board of Radiology certification exams. LLMs have the unique advantage of providing immediate and context-specific explanations of medical physics concepts and are efficient at retrieving information from a wide range of topics. In busy clinical environments, where mentors may sometimes be occupied with clinical duties, LLMs can serve as a useful tool to address easy questions and have trainees pay more attention to more complex problems. This leads to more focused and productive discussions with mentors. In summary, LLMs can bridge gaps in traditional learning settings and support self-directed study in radiation oncology physics.
@@ 5 Conclusion
We evaluated recently released LLMs using a method that randomly shuffled the answer options of radiation oncology physics questions. Our results demonstrated that these models achieved expert-level performance on these questions, with some even surpassing human experts with a majority vote. However, when the correct answer options were replaced with ”None of the above,” all models exhibited a steep decline in performance, suggesting room for improvement. Employing the technique of explain-first and step-by-step instruction enhanced the reasoning abilities of LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet. The performance of these recently released LLMs exhibited great potential for assisting in radiation oncology physics education and training.